ELIMINATE t FROM THE EQUATION x=3+t^3,y=2+1/t

Answer by Alan3354(69443)

ELIMINATE t FROM THE EQUATION x=3+t^3,y=2+1/t
-----
There are 2 equations.
x=3+t^3
y=2+1/t --> t = 1/(y-2)
-----
x = 3 + 1/(y-2)^3 *** No t.